### v2

# Tile the symmetric group by transpositions

Abstract
For nonempty subsets $X$ and $Y$ of a group $G$, we say that $(X,Y)$ is a tiling of $G$ if every element of $G$ can be uniquely expressed as $xy$ for some $x\in X$ and $y\in Y$. In 1966, Rothaus and Thompson studied whether the symmetric group $S_{n}$ with $n\geq 3$ admits a tiling $(T_{n},Y)$, where $T_{n}$ consists of the identity and all the transpositions in $S_{n}$. They showed that no such tiling exists if 1+n(n1)/21+n(n-1)/2 is divisible by a prime number at least $\sqrt{n}+2$. In this paper, we establish a new necessary condition for the existence of such a tiling: the subset $Y$ must be partition-transitive with respect to certain partitions of $n$. This generalizes the result of Rothaus and Thompson, as well as a result of Nomura in 1985. We also study whether $S_{n}$ can be tiled by the set $T_{n}^{*}$ of all the transpositions, which finally leads us to conjecture that neither $T_{n}$ nor $T_{n}^{*}$ tiles $S_{n}$ for any $n\geq 4$.
Key words: tiling; symmetric group; representation; partition-transitivity; perfect code; Cayley graph
AMS subject classification (2020): 20C30, 05E10, 05C25
1 Introduction
Let $G$ be a group. A pair $(X,Y)$ of subsets is called a tiling of $G$ if every element of $G$ can be written as $xy$ for some unique $x\in X$ and $y\in Y$. We say that $X$ left-tiles $G$ if $G$ has a tiling $(X,Y)$ for some subset $Y$. Similarly, $Y$ right-tiles $G$ if $G$ has a tiling $(X,Y)$ for some subset $X$. A subset $S$ is said to tile $G$ if it either left-tiles or right-tiles $G$. Note that, if $S$ is inverse-closed, then $S$ left-tiles $G$ if and only if $S$ right-tiles $G$.
For subsets $X$ and $Y$ of $G$, it is clear that $(X,Y)$ is a tiling of $G$ if and only if $(xX,Yy)$ is a tiling of $G$ for any $x,y\in G$. Therefore, when studying the tilings $(X,Y)$, one usually assumes that both $X$ and $Y$ contain the identity $\mathrm{id}$ of $G$. We refer to such a tiling as normalized. (In some literature, what we call “normalized tilings” are simply referred to as “tilings”.)
Tilings of groups, also referred to as factorizations in the literature (see [30] for example), are closely related to perfect codes in Cayley graphs (see Subsection 1.2). First studied by Hajós [17] in his proof of a well-known conjecture of Minkowski, tilings of abelian groups have received extensive attention [30]. However, much less is known about tilings of nonabelian groups. In 1960s, Rothaus and Thompson [27] considered the possible tilings of the symmetric group $S_{n}$ by the subset
$T_{n}:=\{\mathrm{id}\}\cup\{(i,j)\mid 1\leq i<j\leq n\}$
consisting of the identity and all the transpositions in $S_{n}$. They proved the following result.
Theorem 1.1 (Rothaus–Thompson [27]).
If 1+n(n1)/21+n(n-1)/2 has a prime factor $p\geq\sqrt{n}+2$, then $T_{n}$ does not tile $S_{n}$.
The approach of Rothaus and Thompson [27] to prove the above theorem relies on an ingenious construction of some permutation representation of $S_{n}$, where the condition “1+n(n1)/21+n(n-1)/2 has a prime divisor $p\geq\sqrt{n}+2$” plays an important role. However, following this approach, it seems hard to make further improvements. In fact, necessary conditions for $S_{n}$ to be tiled by $T_{n}$ have been investigated in only a few papers [11] and [26]. Notably, Nomura [26] in 1985 established an interesting result concerning the multiple transitivity of $Y$ in any tiling $(T_{n},Y)$ of $S_{n}$. A subset $Y$ of $S_{n}$ is said to be $k$-transitive if there exists a positive integer $r$ such that, for every pair of $k$-tuples $\alpha$ and $\beta$ of distinct elements in $\{1,\dots,n\}$, there are exactly $r$ permutations in $Y$ mapping $\alpha$ to $\beta$. The main result of [26] is as follows.
Theorem 1.2 (Nomura [26]).
If $(T_{n},Y)$ is a tiling of $S_{n}$, then $Y$ is $k$-transitive for each positive integer $k<n/2$.
In the study of the multiple transitivity and homogeneity for groups and sets of permutations, Martin and Sagan [25] made a remarkable generalization of the notion of transitivity which is closely related to the representation theory of $S_{n}$. Recall that a sequence $(\lambda_{1},\ldots,\lambda_{\ell})$ of integers is called a partition of $n$, denoted $(\lambda_{1},\ldots,\lambda_{\ell})\vdash n$, if $\lambda_{1}\geq\dots\geq\lambda_{\ell}>0$ and $\lambda_{1}+\cdots+\lambda_{\ell}=n$. For convenience, we sometimes treat the sequence $(\lambda_{1},\ldots,\lambda_{\ell})$ as a multiset and use powers to indicate the multiplicities; for example, $(2,1^{3}):=(2,1,1,1)$. An ordered set partition of $\Omega:=\{1,\ldots,n\}$ is a tuple $P=(P_{1},\ldots,P_{\ell})$ of pairwise disjoint subsets of $\Omega$ such that $|P_{1}|\geq\dots\geq|P_{\ell}|>0$ and $P_{1}\cup\dots\cup P_{\ell}=\Omega$. The integer partition $(|P_{1}|,\cdots,|P_{\ell}|)\vdash n$ is called the shape of the ordered set partition $P$.
Definition 1.3 (Martin–Sagan [25]).
Let $\lambda$ be a partition of a positive integer $n$. A subset $Y$ of $S_{n}$ is said to be $\lambda$-transitive if there exits a positive integer $r$ such that, for each ordered set partitions $P$ and $Q$ of shape $\lambda$, there are exactly $r$ permutations in $Y$ sending $P$ to $Q$.
In the terminology of Definition 1.3, the concept of $k$-transitivity is precisely $(n-k,1^{k})$-transitivity. Similarly, $k$-homogeneity is just (nk,k)(n-k,k)-transitivity. In this paper, we first reveal the close relationship between tilings of $S_{n}$ by $T_{n}$ and the notion of partition-transitivity. This does not only give a necessary condition for a tiling $(T_{n},Y)$ of $S_{n}$ in terms of $\lambda$-transitivity for certain partitions $\lambda$ of $n$ (see Theorem 1.4), but also improves both Theorem 1.1 and Theorem 1.2 (see Corollaries 3.4 and 1.6). We then apply the same approach to the tilings $(T_{n}^{*},Y)$ of $S_{n}$, where
$T_{n}^{*}:=T_{n}\setminus\{\mathrm{id}\}=\{(i,j)\mid 1\leq i<j\leq n\}$
is the set of all the transpositions in $S_{n}$. Based on these, we pose a conjecture with some evidence that neither $T_{n}$ nor $T_{n}^{*}$ tiles $S_{n}$ for any $n\geq 4$ (see Section 4).
Main results of the paper
For a partition $\lambda=(\lambda_{1},\dots,\lambda_{\ell})$ of a positive integer $n$, the Young diagram $D(\lambda)$ of $\lambda$ is an array of $n$ boxes having $\ell$ left-justified rows with row $i$ containing exactly $\lambda_{i}$ boxes for $i\in\{1,\dots,\ell\}$. In $D(\lambda)$, the content $\xi(x)$ of a box $x$ in the $i$-th row and $j$-th coloumn is defined by $\xi(x):=j-i$. Our main result is the following theorem and its corollaries.
Theorem 1.4.
If $(T_{n},Y)$ is a tiling of $S_{n}$, then for each $\lambda\vdash n$ with $\sum_{x\in D(\lambda)}\xi(x)\geq 0$, the set $Y$ is $\lambda$-transitive.
Remark.
One can readily verify that, for $\lambda=(\lambda_{1},\dots,\lambda_{\ell})\vdash n$, the content sum
$\sum_{x\in D(\lambda)}\xi(x)=\sum_{i=1}^{\ell}\frac{\lambda_{i}(\lambda_{i}-2i+1)}{2}.$
Thus, $\sum_{x\in D(\lambda)}\xi(x)\geq 0$ if and only if $\sum_{i=1}^{\ell}\lambda_{i}(\lambda_{i}-2i+1)\geq 0$.
Corollary 1.5.
If $T_{n}$ tiles $S_{n}$, then for each partition $\lambda=(\lambda_{1},\dots,\lambda_{\ell})$ of $n$ such that $\sum_{i=1}^{\ell}\lambda_{i}(\lambda_{i}-2i+1)\geq 0$, the integer 1+n(n1)/21+n(n-1)/2 divides $\lambda_{1}!\cdots\lambda_{\ell}!$.
Theorem 1.4 also enables us to obtain the following conclusion that slightly strengthens Nomura’s result (Theorem 1.2).
Corollary 1.6.
If $(T_{n},Y)$ is a tiling of $S_{n}$, then $Y$ is $k$-transitive for each positive integer $k\leq n/2$.
As another application of Theorem 1.4, we are able to slightly weaken the condition $p\geq\sqrt{n}+2$ in Rothaus and Thompson’s result (Theorem 1.1) to $p\geq\sqrt{n}+1$. This leads to a slightly strengthened version of their result in Corollary 3.4.
In [27], Rothaus and Thompson briefly remarked that, if n(n1)/2n(n-1)/2 is divisible by a prime $p\geq\sqrt{n}+2$, then $T_{n}^{*}$ does not tile $S_{n}$. Inspired by this, we establish several results on the tilings of $S_{n}$ by $T_{n}^{*}$ that are parallel to those by $T_{n}$.
Theorem 1.7.
Suppose that $(T_{n}^{*},Y)$ is a tiling of $S_{n}$. Then for each partition $\lambda$ of $n$ with $\sum_{x\in D(\lambda)}\xi(x)>0$, the set $Y$ is $\lambda$-transitive. In particular, $Y$ is $k$-transitive for each positive integer $k\leq(n-2)/2$.
Corollary 1.8.
If $T_{n}^{*}$ tiles $S_{n}$, then for each partition $\lambda=(\lambda_{1},\dots,\lambda_{\ell})$ of $n$ such that $\sum_{i=1}^{\ell}\lambda_{i}(\lambda_{i}-2i+1)>0$, the integer n(n1)/2n(n-1)/2 divides $\lambda_{1}!\cdots\lambda_{\ell}!$.
Based on the above results, we believe that neither $T_{n}$ nor $T_{n}^{*}$ tiles $S_{n}$ for any $n\geq 4$ (see Conjecture 4.1), and various attempts are made in Section 4 towards this conjecture.
More background
The problem whether $T_{n}$ tiles $S_{n}$ is a long-standing open problem. In [29], it is referred to as “the $S_{n}$ problem”. Rothaus and Thompson [27] used the terminology “divide” instead of “tile”. Diaconis [6, Pages 45–46] linked the Rothaus–Thompson approach [27] to the study of random walks on $S_{n}$ and gave the coding-theoretic background of the $S_{n}$ problem, as illustrated below.
In a simple graph $\Gamma$, a perfect code [1] is a set $C$ of vertices such that every vertex of $\Gamma$ is at distance at most one to exactly one vertex in $C$. Given a group $G$ and a subset $S$ of $G$, the Cayley digraph $\mathrm{Cay}(G,S)$ is the digraph with vertex set $G$ and arcs $(g,h)$ whenever $hg^{-1}\in S$. Clearly, $\mathrm{Cay}(G,S)$ is undirected if and only $S$ is inverse-closed, and $\mathrm{Cay}(G,S)$ is loop-free if and only if $\mathrm{id}\notin S$. It is readily seen that a subset $Y$ of $G$ is a perfect code in $\mathrm{Cay}(G,S)$ if and only if every element of $G$ can be written uniquely as a product $xy$ with $x\in S\cup\{\mathrm{id}\}$ and $y\in Y$, or equivalently, if and only if $(S\cup\{\mathrm{id}\},Y)$ is a tiling of $G$. Therefore, the $S_{n}$ problem is precisely concerned with the existence of perfect codes in $\mathrm{Cay}(S_{n},T_{n}^{*})$.
Similarly, a subset $C$ of vertices is said to be a total perfect code [33] in a simple graph $\Gamma$ if every vertex of $\Gamma$ has exactly one neighbor in $C$. Hence, $(T_{n}^{*},Y)$ is a tiling of $S_{n}$ if and only if $Y$ is a total perfect code in $\mathrm{Cay}(S_{n},T_{n}^{*})$. In graph theory, a perfect code is also called an efficient dominating set [4] or independent perfect dominating set [24], and a total perfect code is called an efficient open dominating set [18].
The problem of tiling $S_{n}$ by certain subsets of $T_{n}$ is considered in [7]. As a natural generalization of tilings, an $r$-tiling of a group $G$ is a pair $(X,Y)$ of subsets of $G$ such that every element of $G$ can be written precisely in $r$ ways as $xy$ for some $x\in X$ and $y\in Y$. $r$-Tilings $(X,Y)$ of $S_{n}$ and $\mathrm{SL}_{2}(q)$ such that $X$ is closed under conjugation are studied in [12, 13, 16, 31].
Structure of the paper
After this introduction, we assemble in Section 2 preliminary results on the representations of symmetric groups. Section 3 starts with some technical lemmas for tilings of $S_{n}$ in the language of group algebra and is followed by two more subsections. We tie together all the preparation to prove Theorem 1.4 in Subsection 3.2. Then in Subsection 3.3 some corollaries and related results of Theorem 1.4 are obtained. Finally, we pose the conjecture on the tilings of $S_{n}$ by transpositions, with various attempts, in Section 4.
2 Representation-theoretic results
In this section, we collect some notions and results from representation theory that will be needed in the paper.
Representations of finite groups
Let $\mathbb{C}G$ be the group algebra of a finite group $G$ over $\mathbb{C}$, and identify the representations of $G$ (over $\mathbb{C}$) as $\mathbb{C}G$-modules. A submodule of the left regular representation of $G$ is exactly a left ideal of $\mathbb{C}G$, and an irreducible $\mathbb{C}$-linear representation of $G$ is equivalent to a minimal left ideal of $\mathbb{C}G$. By Maschke’s theorem, each left ideal of $\mathbb{C}G$ is a direct sum of minimal left ideals of $\mathbb{C}G$.
For a character $\chi$ of a representation of $G$, we naturally ($\mathbb{C}$-linearly) extend it to $\chi\colon\mathbb{C}G\rightarrow\mathbb{C}$, and denote
$c_{\chi}:=\dfrac{\chi(\mathrm{id})}{|G|}\sum\limits_{g\in G}\chi(g^{-1})g\ \text{ and }\ I_{\chi}:=(\mathbb{C}G)c_{\chi}.$ (1)
The results in the following lemma are well known and can be read off from [3, §33].
Lemma 2.1.
Let $\chi_{1},\dots,\chi_{r}$ be a complete set of irreducible characters of $G$. Then the following statements hold:
  1. (a)
    $c_{\chi_{1}},\dots,c_{\chi_{r}}$ form a $\mathbb{C}$-basis for the center of $\mathbb{C}G$.
  2. (b)
    For each $i\in\{1,\dots,r\}$, the element $c_{\chi_{i}}$ is the identity of $I_{\chi_{i}}$ and annihilates $I_{\chi_{j}}$ for all $j\in\{1,\dots,r\}\setminus\{i\}$. In particular, each $c_{\chi_{i}}$ is a central idempotent of $\mathbb{C}G$, and $I_{\chi_{i}}I_{\chi_{j}}=I_{\chi_{j}}I_{\chi_{i}}=0$ for distinct $i$ and $j$ in $\{1,\dots,r\}$.
  3. (c)
    $1=\sum_{i=1}^{r}c_{\chi_{i}}$, and hence $\mathbb{C}G=\bigoplus_{i=1}^{r}I_{\chi_{i}}$ is a decomposition of $\mathbb{C}G$ into simple two-sided ideals with the multiplication by $c_{\chi_{i}}$ being the projection into $I_{\chi_{i}}$.
  4. (d)
    For each $i\in\{1,\dots,r\}$, the two-sided ideal $I_{\chi_{i}}$ is the sum of all the minimal left ideals of $\mathbb{C}G$ corresponding to the character $\chi_{i}$.
Representations of $S_{n}$
Let $\lambda=(\lambda_{1},\ldots,\lambda_{\ell})$ and $\mu=(\mu_{1},\ldots,\mu_{m})$ be partitions of $n$. We say that $\mu$ dominates $\lambda$, written $\mu\unrhd\lambda$, if
$\mu_{1}+\mu_{2}+\cdots+\mu_{i}\geq\lambda_{1}+\lambda_{2}+\cdots+\lambda_{i}$
for each positive integer $i$, where, if $i>\ell$ (respectively, $i>m$), then we take $\lambda_{i}$ (respectively, $\mu_{i}$) to be 0. A Young tableau of shape $\lambda$ (also called a $\lambda$-tableau) is an array obtained by replacing the boxes of the Young diagram of $\lambda$ with the numbers $1,2,\ldots,n$ bijectively. For a $\lambda$-tableau $t$, the subgroup $H(t)$ of $S_{n}$ fixing each row of $t$ is called the horizontal group (row group) of $t$, and the subgroup $V(t)$ of $S_{n}$ fixing each column of $t$ is called the vertical group (column group) of $t$.
Let $\lambda=(\lambda_{1},\ldots,\lambda_{\ell})$ be a partition of $n$. Two $\lambda$-tableaux $s$ and $t$ are said to be row-equivalent, written $s\sim t$, if for each $i\in\{1,2,\ldots,\ell\}$, the sets of numbers in the $i$-th row of $s$ and $t$ are equal. A $\lambda$-tabloid is a row-equivalence class of $\lambda$-tableaux, and the $\mathbb{C}S_{n}$-module $M^{\lambda}$ corresponding to $\lambda$ is the $\mathbb{C}$-vector space with the $\lambda$-tabloids as basis and the natural permutation action of $S_{n}$ on $\lambda$-tabloids. For a $\lambda$-tableau $t$, we denote the tabloid $\{s\,|\,s\sim t\}$ by $[t]$ and call
$e_{t}:=\sum_{\sigma\in V(t)}\mathrm{sgn}(\sigma)\sigma[t]$
a $\lambda$-polytabloid. The Specht module $S^{\lambda}$ is the submodule of $M^{\lambda}$ spanned by all $\lambda$-polytabloids. The results in the following lemma can be found in [28, §2.4 and §2.11] and [20, Corollary 2.2.22].
Lemma 2.2.
The following statements hold:
  1. (a)
    The set of Specht modules $\{S^{\lambda}\}_{\lambda\vdash n}$ forms a complete set of pairwise inequivalent irreducible $\mathbb{C}S_{n}$-modules.
  2. (b)
    For each $\lambda\vdash n$, the $\mathbb{C}S_{n}$-module $M^{\lambda}$ decomposes as
    $M^{\lambda}\,\cong\,\bigoplus_{\mu\,\unrhd\,\lambda}K_{\mu\lambda}S^{\mu},$
    where $K_{\mu\lambda}$ is a nonnegative integer, known as the Kostka number, such that $K_{\lambda\lambda}=1$ and that $K_{\mu\lambda}\geq 1$ if and only if $\mu\unrhd\lambda$.
In what follows, we will write $\chi^{\lambda}$ for the character of the irreducible representation $S^{\lambda}$ and write $\chi^{\lambda}(\mu)$ or $\chi^{\lambda}_{\mu}$ for the character value of $\chi^{\lambda}$ at a conjugacy class with cycle type $\mu\vdash n$. With the notation in (1), we write
$c^{\lambda}:=c_{\chi^{\lambda}}\ \text{ and }\ I^{\lambda}:=I_{\chi^{\lambda}}.$ (2)
For a subset $A$ of $S_{n}$, denote
$\overline{A}:=\sum_{a\in A}a\in\mathbb{C}S_{n}.$
Since there is only one $(n)$-tabloid, the $\mathbb{C}S_{n}$-module $M^{(n)}$ is trivial, and so is $S^{(n)}$ by Lemma 2.2. Therefore,
$c^{(n)}=\frac{1}{n!}\overline{S_{n}}\ \text{ and }\ I^{(n)}=\mathbb{C}\overline{S_{n}}.$ (3)
Let $\mathcal{C}_{\mu}$ denote the conjugacy class in $S_{n}$ with cycle type $\mu$. We will also need the notion of the central character
$\omega^{\lambda}_{\mu}:=\frac{|\mathcal{C}_{\mu}|\chi^{\lambda}_{\mu}}{\chi^{\lambda}(\mathrm{id})},$ (4)
where $\chi^{\lambda}(\mathrm{id})=\chi^{\lambda}_{(1^{n})}$ is the degree of $\chi^{\lambda}$. Recall from Subsection 1.1 the definition of the content $\xi(x)$ of a box $x$ in a Young diagram. Interestingly, the central character $\omega^{\lambda}_{(2,1^{n-2})}$ is equal to the content sum $\sum_{x\in D(\lambda)}\xi(x)$ [15, 19] (see also [2]) and is monotonic with respect to the dominance order [5, Lemma 10]. We record these results in the next lemma.
Lemma 2.3.
Let $\lambda$ and $\mu$ be partitions of $n$. Then the following statements hold:
  1. (a)
    $\omega^{\lambda}_{(2,1^{n-2})}=\sum_{x\in D(\lambda)}\xi(x)$.
  2. (b)
    If $\mu\unrhd\lambda$ and $\mu\neq\lambda$, then $\omega^{\mu}_{(2,1^{n-2})}>\omega^{\lambda}_{(2,1^{n-2})}$.
Both $M^{\lambda}$ and $S^{\lambda}$ have their left ideal avatars in the group algebra $\mathbb{C}S_{n}$, which we now describe. Let $\lambda\vdash n$, and let $t$ be a $\lambda$-tableau. For a subgroup $H$ of $S_{n}$, the mapping $w\overline{H}\mapsto wH$ (this mapping is the $\mathbb{C}$-linear span of the mapping $g\overline{H}\mapsto gH$ for each $g\in S_{n}$) is a $\mathbb{C}S_{n}$-module isomorphism from $(\mathbb{C}S_{n})\overline{H}$ to the permutation representation of $S_{n}$ by left multiplication on the set $S_{n}/H$ of the left cosets of $H$. Hence, $(\mathbb{C}S_{n})\overline{H(t)}$ is isomorphic to $M^{\lambda}$ by [28, Theorem 2.1.12]. Denote
$E_{t}:=\sum_{\pi\in V(t)}\sum_{\rho\in H(t)}\mathrm{sgn}(\pi)\pi\rho\ \text{ and }\ L^{\lambda}:=(\mathbb{C}S_{n})E_{t}.$
It turns out that $L^{\lambda}$ is an avatar of $S^{\lambda}$ in $\mathbb{C}S_{n}$ (see for instance [20, Lemma 7.1.4]). Therefore, we have the following lemma.
Lemma 2.4.
Let $\lambda\vdash n$, and let $t$ be a $\lambda$-tableau. Then there hold the following isomorphisms of $\mathbb{C}S_{n}$-modules:
  1. (a)
    The left ideal $(\mathbb{C}S_{n})\overline{H(t)}$ of $\mathbb{C}S_{n}$ is isomorphic to $M^{\lambda}$.
  2. (b)
    The left ideal $L^{\lambda}$ of $\mathbb{C}S_{n}$ is isomorphic to $S^{\lambda}$.
3 $\lambda$-transitivity and tilings of $S_{n}$ by transpositions
In this section, we prove Theorem 1.4–Corollary 1.8, as well as a slightly strengthened version (see Corollary 3.4) of Rothaus–Thompson’s result (Theorem 1.1).
Technical lemmas
By definition, $(X,Y)$ is a tiling of a group $G$ if and only if for each $g\in G$ there exists a unique $(x,y)\in X\times Y$ satisfying $g=xy$, which is equivalent to the equation $\overline{G}=\overline{X}\,\overline{Y}$ in $\mathbb{C}G$. For a subset $X$ that is closed under conjugation in $G$, the pair $(X,Y)$ is a tiling of $G$ if and only if $(X,gYh)$ is a tiling of $G$ for any $g,h\in G$. Moreover, if $X$ is closed under conjugation in $G$, then $\overline{X}$ is central in $\mathbb{C}G$ and so $\overline{X}\,\overline{Y}=\overline{Y}\,\overline{X}$. In this case, if $X$ is inverse-closed in addition, then $(X,Y)$ being a tiling of $G$ is equivalent to $(X,Y^{-1})$ being a tiling of $G$, where $Y^{-1}:=\{y^{-1}\,|\,y\in Y\}$. We summarize these observations in the following lemma.
Lemma 3.1.
Let $X$ and $Y$ be subsets of a group $G$. Then the following statements hold:
  1. (a)
    $(X,Y)$ is a tiling of $G$ if and only if $\overline{G}=\overline{X}\,\overline{Y}$.
  2. (b)
    If $(X,Y)$ is a tiling of $G$ such that $X$ is closed under conjugation in $G$, then $(X,gYh)$ is also a tiling of $G$ for all $g,h\in G$.
  3. (c)
    If $(X,Y)$ is a tiling of $G$ such that $X$ is closed under inversion and conjugation in $G$, then $(X,Y^{-1})$ is also a tiling of $G$.
Since $T_{n}$ is closed under conjugation in $S_{n}$, the element $\overline{T_{n}}$ of $\mathbb{C}S_{n}$ is central. Recall the definition of $c^{\lambda}$ and $I^{\lambda}$ in (2), where $\lambda$ is a partition of $n$, and recall the definition of the central character in (4). The next lemma shows that the multiplication by $\overline{T_{n}}$ restricted on $I^{\lambda}$ is the scalar multiplication of $1+\omega^{\lambda}_{(2,1^{n-2})}$.
Lemma 3.2.
Let $\lambda\vdash n$. Then $\overline{T_{n}}\,u=u\,\overline{T_{n}}=\big(1+\omega^{\lambda}_{(2,1^{n-2})}\big)u$ for all $u\in I^{\lambda}$.
Proof.  Consider the minimal left ideal $L^{\lambda}$ (see Lemmas 2.2 and 2.4) and the mapping
$\varphi\colon L^{\lambda}\rightarrow L^{\lambda},\ \ u\mapsto\overline{T_{n}}\,u.$
Since $\overline{T_{n}}$ lies in the center of $\mathbb{C}S_{n}$, then by Schur’s lemma, there exists some constant $d\in\mathbb{C}$ such that $\varphi(u)=du$ for all $u\in L^{\lambda}$. Hence we have
$\chi^{\lambda}(\mathrm{id})d=\mathrm{Tr}(\varphi)=\chi^{\lambda}(\overline{T_{n}})=\chi^{\lambda}(\mathrm{id})+|\mathcal{C}_{(2,1^{n-2})}|\chi^{\lambda}_{(2,1^{n-2})}=\chi^{\lambda}(\mathrm{id})+\chi^{\lambda}(\mathrm{id})\omega^{\lambda}_{(2,1^{n-2})},$
where $\mathrm{Tr}(\varphi)$ is the trace of $\varphi$. This leads to $d=1+\omega^{\lambda}_{(2,1^{n-2})}$. In other words, $\overline{T_{n}}\,u=\big(1+\omega^{\lambda}_{(2,1^{n-2})}\big)u$ for all $u\in L^{\lambda}$. Note from Lemma 2.1 that $I^{\lambda}$ is the sum of all the minimal left ideals of $\mathbb{C}G$ corresponding to the character $\chi^{\lambda}$. We conclude that $\overline{T_{n}}\,u=\big(1+\omega^{\lambda}_{(2,1^{n-2})}\big)u$ for all $u\in I^{\lambda}$, as required. ∎
Based on the above discussion, we obtain the following necessary condition for $(T_{n},Y)$ to be a tiling of $S_{n}$.
Lemma 3.3.
Suppose that $(T_{n},Y)$ is a tiling of $S_{n}$ with $n\geq 3$. Then $c^{\lambda}\overline{Y}=0$ for each $\lambda\vdash n$ such that $\lambda\neq(n)$ and $\omega^{\lambda}_{(2,1^{n-2})}\neq-1$, or equivalently, $\overline{Y}\in I^{(n)}\boldsymbol{\oplus}\bigg(\bigoplus_{\lambda\in\Lambda}I^{\lambda}\bigg)$, where $\Lambda:=\{\lambda\vdash n\mid\omega^{\lambda}_{(2,1^{n-2})}=-1\}$.
Proof.  Let $\lambda\vdash n$ such that $\lambda\neq(n)$ and $\omega^{\lambda}_{(2,1^{n-2})}\neq-1$. Note from (3) that $I^{(n)}=\mathbb{C}\overline{S_{n}}$. Then $c^{\lambda}\overline{S_{n}}=0$ by Lemma 2.1, and $c^{\lambda}\,\overline{T_{n}}=\big(1+\omega^{\lambda}_{(2,1^{n-2})}\big)c^{\lambda}$ by Lemma 3.2. Since $(T_{n},Y)$ is a tiling of $S_{n}$, it follows from Lemma 3.1 that $\overline{T_{n}}\,\overline{Y}=\overline{S_{n}}$, which yields
$0=c^{\lambda}\overline{S_{n}}=c^{\lambda}\overline{T_{n}}\,\overline{Y}=\big(1+\omega^{\lambda}_{(2,1^{n-2})}\big)(c^{\lambda}\overline{Y}).$
Thus, $c^{\lambda}\overline{Y}=0$, as required. To see the equivalent conclusion, notice Lemma 2.1 and that $(n)\notin\Lambda$ as $\omega^{(n)}_{(2,1^{n-2})}=n(n-1)/2$ by Lemma 2.3. ∎
Proof of Theorem 1.4 and Corollary 1.5
With the preparation so far, we now embark on the proof of Theorem 1.4 and Corollary 1.5. They hold trivially for $n=2$. Hence we assume $n\geq 3$ in the following. Let $(T_{n},Y)$ be a tiling of $S_{n}$, and let $\lambda$ be a partition of $n$ such that $\sum_{x\in D(\lambda)}\xi(x)\geq 0$.
Recall from Subsection 2.2 the definition of the $\mathbb{C}S_{n}$-module $M^{\lambda}$. By Lemmas 2.2 and 2.4, there exists a $\mathbb{C}S_{n}$-isomorphism $\rho$ from $M^{\lambda}$ to $\bigoplus_{\mu\,\unrhd\,\lambda}K_{\mu\lambda}L^{\mu}$. Let $[s]$ and $[t]$ be arbitrary $\lambda$-tabloids. Then $h[s]=[t]$ for some $h\in S_{n}$, and so we derive from (3) that
$c^{(n)}\rho([s]-[t])$ $=c^{(n)}\big(\rho([s])-h\rho([s])\big)$
$=\frac{1}{n!}\,\overline{S_{n}}\big(\rho([s])-h\rho([s])\big)=\frac{1}{n!}\,\overline{S_{n}}\,\rho([s])-\frac{1}{n!}\,\overline{S_{n}}\,\rho([s])=0.$
This together with Lemma 2.1 implies that
$\rho([s]-[t])\in\bigoplus_{\begin{subarray}{c}\mu\,\unrhd\,\lambda\\ \mu\neq(n)\end{subarray}}K_{\mu\lambda}L^{\mu}.$ (5)
For each $\mu\unrhd\lambda$ with $\mu\neq(n)$, we have by Lemma 2.3 that $\omega^{\mu}_{(2,1^{n-2})}\geq\omega^{\lambda}_{(2,1^{n-2})}\geq 0$ (in particular, $\omega^{\mu}_{(2,1^{n-2})}\neq-1$), and thus $\overline{Y}L^{\mu}=0$ by Lemma 2.1. It then follows from (5) that $\overline{Y}\rho([s]-[t])=0$, which is equivalent to $\overline{Y}([s]-[t])=0$. According to Lemma 3.1, $(T_{n},Y^{-1})$ is also a tiling of $S_{n}$, whence we have $\overline{Y^{-1}}([s]-[t])=0$ in addition. Therefore,
$\overline{Y}\,[s]=\overline{Y}\,[t]\ \text{ and }\ \overline{Y^{-1}}\,[s]=\overline{Y^{-1}}\,[t].$ (6)
Now fix a $\lambda$-tabloid $s_{0}]$, and let $r$ be the numbers of elements $y$ in $Y$ such that $y^{-1}[s_{0}]=[s_{0}]$. For arbitrary $\lambda$-tabloids $s_{1}$] and $s_{2}$], it follows from (6) that $\overline{Y}\,[s_{1}]=\overline{Y}\,[s_{0}]$ and $\overline{Y^{-1}}\,[s_{2}]=\overline{Y^{-1}}\,[s_{0}]$. Comparing the coefficients of $s_{0}]$ in both sides of $\overline{Y^{-1}}\,[s_{2}]=\overline{Y^{-1}}\,[s_{0}]$, we obtain
$|\{y\in Y\mid y^{-1}[s_{2}]=[s_{0}]\}|=|\{y\in Y\mid y^{-1}[s_{0}]=[s_{0}]\}|=r.$
Then comparing the coefficients of $s_{2}]$ in both sides of $\overline{Y}\,[s_{1}]=\overline{Y}\,[s_{0}]$, we infer that
$|\{y\in Y\mid y[s_{1}]=[s_{2}]\}|$ $=|\{y\in Y\mid y[s_{0}]=[s_{2}]\}|$
$=|\{y\in Y\mid y^{-1}[s_{2}]=[s_{0}]\}|=r.$
By the definition of $\lambda$-transitivity and $M^{\lambda}$, this means that for each ordered set partitions $P$ and $Q$ of shape $\lambda$, there are exactly $r$ permutations in $Y$ sending $P$ to $Q$. Write $\lambda=(\lambda_{1},\ldots,\lambda_{\ell})$. Then there are precisely $n!/\lambda_{1}!\cdots\lambda_{\ell}!$ ordered set partitions of shape $\lambda$, and so $r=|Y|/(n!/\lambda_{1}!\cdots\lambda_{\ell}!)>0$. Thus, according to the definition of $\lambda$-transitivity, $Y$ is $\lambda$-transitive, as Theorem 1.4 asserts. Moreover, since $r$ is an integer, $n!/\lambda_{1}!\cdots\lambda_{\ell}!$ divides $|Y|=|S_{n}|/|T_{n}|=n!/|T_{n}|$. It follows that $1+n(n-1)/2=|T_{n}|$ divides $\lambda_{1}!\cdots\lambda_{\ell}!$, which together with the Remark after Theorem 1.4 implies Corollary 1.5. ∎
Other results
Corollary 3.4 (A strengthened version of [27, Theorem]).
If 1+n(n1)/21+n(n-1)/2 has a prime divisor $p\geq\sqrt{n}+1$, then $T_{n}$ does not tile $S_{n}$.
Proof.  Suppose on the contrary that there exists some tiling $(T_{n},Y)$ of $S_{n}$. Write n=(p1)q+rn=(p-1)q+r with nonnegative integers $q$ and $r$ such that $r\leq p-2$. Consider the partition $\lambda=(\lambda_{1},\ldots,\lambda_{\ell})=((p-1)^{q},r)\vdash n$. Then
$\sum_{i=1}^{\ell}\lambda_{i}(\lambda_{i}-2i+1)$ $=\sum_{i=1}^{q}(p-1)(p-2i)+r(r-2q-1)$
$=(p-1)(p-1-q)q+r(r-2q-1).$ (7)
By Corollary 1.5, it suffices to prove that (7) is nonnegative. Since $n=(p-1)q+r\geq(p-1)q$ and $p\geq\sqrt{n}+1$, we have
$q\leq\frac{n}{p-1}\leq\frac{n}{\sqrt{n}}=\sqrt{n}\leq p-1.$
If q=p1q=p-1, then $\sqrt{n}=p-1=q$ and hence r=n(p1)q=0r=n-(p-1)q=0, which yields that (7) equals 0. Now assume that $q\leq p-2$. Then since the minimum of r(r2q1)r(r-2q-1) over integers $r$ is (q+1)q-(q+1)q, it follows that (7) is at least
$(p-1)(p-1-q)q-(q+1)q\geq\big((q+2)-1\big)\big((q+2)-1-q\big)q-(q+1)q=0.$
This completes the proof. ∎
Next we give a proof of Corollary 1.6:
Consider the hook-like partition $\lambda=(n-k,1^{k})\vdash n$. It satisfies $\sum_{x\in D(\lambda)}\xi(x)\geq 0$ if and only if $n-k\geq k+1$. If $k<n/2$, then $Y$ is $\lambda$-transitive by Theorem 1.4. Thus we only need to deal with the case when $n=2k\geq 4$ is even. As in the proof of Theorem 1.4, it suffices to show that in the decomposition $M^{\lambda}\cong\bigoplus_{\mu\,\unrhd\,\lambda}K_{\mu\lambda}L^{\mu}$ we have $\omega^{\mu}_{(2,1^{n-2})}\neq-1$ whenever $\mu\unrhd\lambda$. Clearly, $\omega^{\lambda}_{(2,1^{n-2})}=-k\neq-1$. If $\mu\unrhd\lambda$ and $\mu\neq\lambda$, then $\mu\unrhd(k,2,1^{k-2})$ and hence we deduce from Lemma 2.3 that $\omega^{\mu}_{(2,1^{n-2})}\geq\omega^{(k,2,1^{k-2})}_{(2,1^{n-2})}=0$. ∎
Now consider the tilings of $S_{n}$ by $T_{n}^{*}$. Through a very similar argument as the above proof of Theorem 1.4 and its corollaries, we give a proof of Theorem 1.7 and Corollary 1.8:
By Lemma 3.2, $\overline{T_{n}^{*}}\,u=u\,\overline{T_{n}^{*}}=\omega^{\lambda}_{(2,1^{n-2})}u$ for all $\lambda\vdash n$ and $u\in I^{\lambda}$. This implies that, similarly to Lemma 3.3, if $(T_{n}^{*},Y)$ is a tiling of $S_{n}$ with $n\geq 3$, then $c^{\lambda}\overline{Y}=0$ for each $\lambda\vdash n$ with $\lambda\neq(n)$ and $\omega^{\lambda}_{(2,1^{n-2})}\neq 0$. Then by the argument in the proof of Theorem 1.4 and Corollary 1.5, one proves Theorem 1.7 and Corollary 1.8. ∎
4 Conjectures and discussions
In the proceeding sections we already see that, in a tiling $(T_{n},Y)$ of $S_{n}$, the subset $Y$ is restricted in size, transitivity, scope where $\overline{Y}$ belongs to (see Lemma 3.3). Some other requirements will also be revealed later in this section. Based on this information, we believe that such a subset $Y$ can hardly exist, and thus the following conjecture seems reasonable.
Conjecture 4.1.
For $n\geq 4$, neither $T_{n}$ nor $T_{n}^{*}$ tiles $S_{n}$.
It can be directly verified for $n=4$ that neither $T_{n}$ nor $T_{n}^{*}$ tiles $S_{n}$. Thus, assume
$n\geq 5$
in the following discussion. By Corollary 1.6 and Theorem 1.7, if either $T_{n}$ or $T_{n}^{*}$ tiles $S_{n}$, then there exists some $\lfloor(n-2)/2\rfloor$-transitive subset $Y$ of $S_{n}$ such that $Y$ is not $A_{n}$ or $S_{n}$. Given any positive integers $\lambda_{2}\geq\dots\geq\lambda_{\ell}$, Martin and Sagan [25, Section 6] constructed infinitely many values for $\lambda_{1}\geq\lambda_{2}$ such that there exists a $\lambda$-transitive subset of $S_{n}$ with $\lambda=(\lambda_{1},\lambda_{2},\ldots,\lambda_{\ell})$. For each positive integer $k$, by taking $\ell=k+1$ and $\lambda_{2}=\cdots=\lambda_{\ell}=1$ we see that there exist $k$-transitive subsets of $S_{n}$ for infinitely many values of $n$. Since this only gives examples of $k$-transitive subsets of $S_{n}$ with $k/n\to 0$ whereas $\lfloor(n-2)/2\rfloor/n\to 1/2$, we pose the following conjecture, which will imply Conjecture 4.1 for sufficiently large $n$.
Conjecture 4.2.
For sufficiently large $n$, there is no $\lfloor(n-2)/2\rfloor$-transitive subset of $S_{n}$ other than $A_{n}$ and $S_{n}$.
In the rest of this section we focus on Conjecture 4.1, providing more insights into this conjecture. Denote $T_{n}^{2}:=\{t_{1}t_{2}\mid t_{1},t_{2}\in T_{n}\}$ and $(T_{n}^{*})^{2}:=\{t_{1}t_{2}\mid t_{1},t_{2}\in T_{n}^{*}\}$. Then
$T_{n}^{2}\setminus\{\mathrm{id}\}=\mathcal{C}_{(2,1^{n-2})}\cup\mathcal{C}_{(3,1^{n-3})}\cup\mathcal{C}_{(2^{2},1^{n-4})}\ \text{ and }\ (T_{n}^{*})^{2}\setminus\{\mathrm{id}\}=\mathcal{C}_{(3,1^{n-3})}\cup\mathcal{C}_{(2^{2},1^{n-4})}.$
As usual, $\alpha(\Gamma)$ denotes the independence number of a graph $\Gamma$, namely, the maximum size of an independent set (coclique) of $\Gamma$. Let us start with the following observation.
Lemma 4.3.
Let $\Sigma_{n}:=\mathrm{Cay}(S_{n},T_{n}^{2}\setminus\{\mathrm{id}\})$ and $\Sigma_{n}^{*}:=\mathrm{Cay}(S_{n},(T_{n}^{*})^{2}\setminus\{\mathrm{id}\})$. Then the following statements hold:
  1. (a)
    $(T_{n},Y)$ is a tiling of $S_{n}$ if and only if $Y$ is an independent set of size n!/(1+n(n1)/2)n!/(1+n(n-1)/2) in $\Sigma_{n}$.
  2. (b)
    $(T_{n}^{*},Y)$ is a tiling of $S_{n}$ if and only if $Y$ is an independent set of size 2(n2)!2(n-2)! in $\Sigma_{n}^{*}$.
  3. (c)
    $T_{n}$ tiles $S_{n}$ if and only if $\alpha(\Sigma_{n})\geq n!/(1+n(n-1)/2)$.
  4. (d)
    $T_{n}$ tiles $S_{n}$ if and only if $\alpha(\Sigma_{n})=n!/(1+n(n-1)/2)$.
  5. (e)
    $T_{n}^{*}$ tiles $S_{n}$ if and only if $\alpha(\Sigma_{n}^{*})\geq 2(n-2)!$.
  6. (f)
    $T_{n}^{*}$ tiles $S_{n}$ if and only if $\alpha(\Sigma_{n})=2(n-2)!$.
Proof.  By the definition of a tiling, $(T_{n},Y)$ is a tiling of $S_{n}$ if and only if $|Y|=n!/|T_{n}|$ and there do not exist distinct $y_{1}$ and $y_{2}$ in $Y$ and distinct $t_{1}$ and $t_{2}$ in $T_{n}$ such that $t_{1}y_{1}=t_{2}y_{2}$. Notice that $|T_{n}|=1+n(n-1)/2$ and
$t_{1}y_{1}=t_{2}y_{2}\;\Leftrightarrow\;y_{1}y_{2}^{-1}=t_{1}^{-1}t_{2}\;\Leftrightarrow\;y_{1}y_{2}^{-1}=t_{1}t_{2}.$ (8)
We conclude that $(T_{n},Y)$ is a tiling of $S_{n}$ if and only if |Y|=n!/(1+n(n1)/2)|Y|=n!/(1+n(n-1)/2) and $Y$ is an independent set of $\mathrm{Cay}(S_{n},T_{n}^{2}\setminus\{\mathrm{id}\})=\Sigma_{n}$. This proves statement (a), and a similar argument gives a proof of statement (b). Note that (8) also shows $T_{n}y_{1}\cap T_{n}y_{2}=\varnothing$ for any distinct $y_{1}$ and $y_{2}$ from an independent set of $\Sigma_{n}$. Hence $\alpha(\Sigma_{n})\leq|S_{n}|/|T_{n}|=n!/(1+n(n-1)/2)$, and statements (c) and (d) follow from (a). Similarly, statements (e) and (f) are consequences of (b). ∎
Fourier transform of Boolean functions on $S_{n}$
For a group $G$ and a complex-valued function $f$ on $G$, the Fourier transform of $f$ is the matrix-valued function $\widehat{f}$ on the set of irreducible representations of $G$ such that
$\widehat{f}(\rho):=\frac{1}{|G|}\sum_{g\in G}f(g)\rho(g)$
for each irreducible representation $\rho$ of $G$. In particular, for a subset $Y$ of $S_{n}$, the Fourier transform of the characteristic function $\mathrm{1}_{Y}$ satisfies
$\widehat{1_{Y}}(\rho)=\frac{1}{n!}\sum_{g\in S_{n}}1_{Y}(g)\rho(g)=\frac{1}{n!}\sum_{g\in Y}\rho(g).$
Note that the only partitions $\lambda\vdash n$ that do not satisfy $\lambda\unrhd(3,1^{n-3})$ are $\lambda=(2^{k},1^{n-2k})$ with nonnegative $k$, which are such that $\sum_{x\in D(\lambda)}\xi(x)<-1$ as $n\geq 5$. If $(T_{n},Y)$ is a tiling of $S_{n}$, then we derive from Lemmas 2.3 and 3.3 that $\widehat{1_{Y}}$ is supported only on irreducible representations $S^{\lambda}$ with $\lambda\unrhd(3,1^{n-3})$. Similarly, if $(T_{n}^{*},Y)$ is a tiling of $S_{n}$, we also have that $\widehat{1_{Y}}$ is supported only on irreducible representations $S^{\lambda}$ with $\lambda\unrhd(3,1^{n-3})$.
As a statement concerning Boolean functions on $S_{n}$ and their Fourier transforms, [8, Theorem 27] would imply Conjecture 4.1. In fact, based on the above conclusion, [8, Theorem 27] asserts that the set $Y$ in any tiling $(T_{n},Y)$ or $(T_{n}^{*},Y)$ is a disjoint union of left cosets of stabilizers in $S_{n}$ of n3n-3 points. Then by Lemma 3.1, the set $Y^{-1}$ in any tiling $(T_{n},Y)$ or $(T_{n}^{*},Y)$ is a disjoint union of left cosets of stabilizers in $S_{n}$ of n3n-3 points. In this case, for any $y\in Y$, there exists a $3$-cycle $(i,j,k)$ in $S_{n}$ such that $y^{-1}(i,j,k)\in Y^{-1}$, which implies that $(i,k,j)y\in Y$, contradicting Lemma 4.3.
Unfortunately, [8, Theorem 27] is found to be false [9, 14].
Eigenvalues of Cayley graphs
Recall from the concept of an $r$-tiling (see Subsection 1.2) that a $1$-tiling of a group $G$ is just a tiling of $G$. The following lemma gives a necessary condition for $r$-tilings in terms of eigenvalues of Cayley digraphs.
Lemma 4.4 ([31, Lemma 1]).
If $(X,Y)$ is an $r$-tiling of a group $G$ for some positive integer $r$ such that $Y\neq G$, then 0 is an eigenvalue of $\mathrm{Cay}(G,X)$.
Based on Lemma 2.1, it is easy to know (refer to [32]) the eigenvalues and eigenvectors of a Cayley digraph $\mathrm{Cay}(G,S)$ with $S$ normal (closed under conjugation) in $G$.
Lemma 4.5.
Let $G$ be a group of order $m$. Then there exist pairwise orthogonal vectors $v_{1},\dots,v_{m}\in\mathbb{C}^{m}$ such that, for each normal subset $S$ of $G$, the adjacency matrix of $\mathrm{Cay}(G,S)$ has eigenvectors $v_{1},\dots,v_{m}\in\mathbb{C}^{m}$ with eigenvalues
$\frac{1}{\chi(\mathrm{id})}\sum_{g\in S}\chi(g),$
where $\chi$ runs over the irreducible characters of $G$.
Then by Lemma 2.3, the eigenvalues of $\mathrm{Cay}(S_{n},T_{n})$ are $1+\sum_{x\in D(\lambda)}\xi(x)$ with partitions $\lambda$ of $n$. This result in conjunction with Lemma 4.4 can be used to show the nonexistence of tilings $(T_{n},Y)$ of $S_{n}$ for some small values of $n$. For example, there is no partition $\lambda\vdash 6$ with content sum $\sum_{x\in D(\lambda)}\xi(x)=-1$, and so 0 is not an eigenvalue of $\mathrm{Cay}(S_{6},T_{6})$, which implies the nonexistence of tilings $(T_{6},Y)$ of $S_{6}$ (note that this nonexistence result cannot be obtained from Lemma 1.5 or 3.4). For a general $n$, however, this approach does not work, as there always exists $\lambda\vdash n$ with $\sum_{x\in D(\lambda)}\xi(x)=-1$ whenever $n\geq 14$. Similarly, since $n$ always has a partition with content sum 0, Lemma 4.4 does not help to exclude the tilings $(T_{n}^{*},Y)$ of $S_{n}$.
Recall from Lemma 4.3 the definitions of $\Sigma_{n}$ and $\Sigma_{n}^{*}$. Inspired by Lemma 4.3, we now turn to the independence numbers of $\Sigma_{n}$ and $\Sigma_{n}^{*}$, which are well known to be related to the eigenvalues of these Cayley graphs, for instance, by the following result [8, Theorem 12].
Lemma 4.6 (Weighted version of Hoffman bound).
Let $\Gamma$ be a graph on $m$ vertices, let $\Gamma_{j}$ be a $d_{j}$-regular spanning subgraph of $\Gamma$ for $j\in\{1,\ldots,r\}$ such that all $\Gamma_{j}$ share a common orthonormal system of eigenvectors $v_{1},\ldots,v_{m}$, and let $\beta_{1},\ldots,\beta_{r}\in\mathbb{R}$. For each $i\in\{1,\ldots,m\}$, let $\ell_{i}:=\sum_{j=1}^{r}\beta_{j}\ell_{ij}$, where $\ell_{ij}$ is the eigenvalue of $v_{i}$ in $\Gamma_{j}$. If $Y$ is an independent set of $\Gamma$, then setting $d:=\sum_{j=1}^{r}\beta_{j}d_{j}$ and $\ell_{\mathrm{min}}:=\min_{i\in\{1,\ldots,m\}}\ell_{i}$, we have
$\frac{|Y|}{m}\leq\frac{-\ell_{\mathrm{min}}}{d-\ell_{\min}};$
moreover, the equality holds only if the characteristic vector $1_{Y}$ of $Y$ is a linear combination of the all-$1$’s vector and those $v_{i}$ with $\ell_{i}=\ell_{\min}$.
By Lemma 4.5, if we take $\Gamma=\Sigma_{n}$ in Lemma 4.6 with
$\Gamma_{1}=\mathrm{Cay}\big(S_{n},\mathcal{C}_{(2,1^{n-2})}\big),\ \ \Gamma_{2}=\mathrm{Cay}\big(S_{n},\mathcal{C}_{(3,1^{n-3})}\big),\ \ \Gamma_{3}=\mathrm{Cay}\big(S_{n},\mathcal{C}_{(2^{2},1^{n-4})}\big),$
then the $\ell_{i}$’s turn out to be
$\ell_{\lambda}:=\beta_{1}\omega^{\lambda}_{(2,1^{n-2})}+\beta_{2}\omega^{\lambda}_{(3,1^{n-3})}+\beta_{3}\omega^{\lambda}_{(2^{2},1^{n-4})}$
with $\lambda$ running over the partitions of $n$. According to [21, Pages 150 and 152],
$\omega^{\lambda}_{(2,1^{n-2})}$ $=\sum_{x\in D(\lambda)}\xi(x),$
$\omega^{\lambda}_{(3,1^{n-3})}$ $=\sum_{x\in D(\lambda)}\xi(x)^{2}-\dfrac{n(n-1)}{2},$
$\omega^{\lambda}_{(2^{2},1^{n-4})}$ $=\dfrac{1}{2}\Big(\sum_{x\in D(\lambda)}\xi(x)\Big)^{2}-\dfrac{3}{2}\sum_{x\in D(\lambda)}\xi(x)^{2}+\dfrac{n(n-1)}{2}.$
Denote $d_{1}:=\big|\mathcal{C}_{(2,1^{n-2})}\big|$, $d_{2}:=\big|\mathcal{C}_{(3,1^{n-3})}\big|$ and $d_{3}:=\big|\mathcal{C}_{(2^{2},1^{n-4})}\big|$. Then
$d_{1}=\frac{n(n-1)}{2},\ \ d_{2}=\frac{n(n-1)(n-2)}{3},\ \ d_{3}=\frac{n(n-1)(n-2)(n-3)}{8}.$
Suppose that $(T_{n},Y)$ is a tiling of $S_{n}$. Then a combination of Lemmas 4.3 and 4.6 gives
$\frac{1}{|T_{n}|}=\frac{|Y|}{n!}\leq\frac{-\ell_{\min}}{\beta_{1}d_{1}+\beta_{2}d_{2}+\beta_{3}d_{3}-\ell_{\min}}$ (9)
for any $\beta_{1},\beta_{2},\beta_{3}\in\mathbb{R}$, where $\ell_{\min}$ is the minimum of $\ell_{\lambda}$ for $\lambda\vdash n$. For $(\beta_{1},\beta_{2},\beta_{3})=(2,3,2)$, the minimum of $\ell_{\min}$ is taken when $\sum_{x\in D(\lambda)}\xi(x)=-1$, and so (9) turns out to be
$\frac{1}{|T_{n}|}\leq\frac{\frac{n(n-1)}{2}+1}{2d_{1}+3d_{2}+2d_{3}+\frac{n(n-1)}{2}+1}=\frac{2}{n^{2}-n+2}.$
However, this is a tight bound for $|Y|/n!$ instead of a contradiction, and the “moreover” part in Lemma 4.6 only yields the conclusion of Lemma 3.3. Similarly, if $(T_{n}^{*},Y)$ is a tiling of $S_{n}$, then taking $(\beta_{1},\beta_{2},\beta_{3})=(0,3,2)$ we obtain $\ell_{\min}$ when $\sum_{x\in D(\lambda)}\xi(x)=0$, which gives the tight bound
$\frac{1}{|T_{n}^{*}|}\leq\frac{\frac{n(n-1)}{2}}{3d_{2}+2d_{3}+\frac{n(n-1)}{2}}=\frac{2}{n^{2}-n}$
for $|Y|/n!$. Currently, we do not know whether it works to prove Conjecture 4.1 by taking certain values of $\beta_{1},\beta_{2},\beta_{3}\in\mathbb{R}$, possibly depending on $n$.
Forbidden intersection problem for permutations
For $i\in\{1,\ldots,n\}$, a subset $Y$ of $S_{n}$ is said to avoid intersection $i$ if the number of fixed points of $y_{1}y_{2}^{-1}$ is not equal to $i$ for any $y_{1},y_{2}\in Y$. By Lemma 4.3, the set $Y$ in any tiling $(T_{n},Y)$ or $(T_{n}^{*},Y)$ of $S_{n}$ avoids intersection n3n-3. Thus, a necessary condition for $T_{n}$ or $T_{n}^{*}$ to tile $S_{n}$ is the existence of a set of size n!/(1+n(n1)/2)n!/(1+n(n-1)/2) or 2(n2)!2(n-2)!, respectively, in $S_{n}$ that avoids intersection n3n-3. The forbidden intersection problem for permutations concerns the maximum size for a subset of $S_{n}$ that avoids intersection $i$, which has attracted considerable attention with breakthroughs achieved recently [10, 22, 23]. However, existing results pertain only to relatively small $i$ (for example, $i=O(\sqrt{n}/\log_{2}n)$ [23, Theorem 2]), while the results we need for Conjecture 4.1 is i=n3i=n-3, as posed in the following question.
Question 4.7.
What is the maximum size for a subset of $S_{n}$ that avoids intersection n3n-3?
Acknowledgements
The authors would like to thank the anonymous referee for helpful comments that have improved the presentation. We also wish to thank Professors William Yong-Chuan Chen and Xin-Gui Fang for their advice and inspiration. Teng Fang was supported by the Scientific Research Foundation for Advanced Talents of Suqian University (no. 2025XRC027). Binzhou Xia acknowledges the support of ARC Discovery Project DP250104965.